OpenStudy (anonymous):

how do i add 3 3/8 and 3 7/8 in fraction form?

OpenStudy (anonymous):

Can you tell me what both of these fractions are when written as improper fractions?

OpenStudy (anonymous):

27/8 and 31/8

OpenStudy (anonymous):

Exactly, now we have a common denominator, we can simply add the numerators and you have your answer. Can you tell me what it is?

OpenStudy (anonymous):

58/8?

OpenStudy (anonymous):

Spot on! Great work!

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

so how is that simplified?

hero (hero):

Why do you need to put it in fraction form? Just add it.

hero (hero):

You have 3 + 3/8 + 3 + 7/8 Put like terms together: 3 + 3 + 3/8 + 7/8 Now you have 6 + 10/8 Which expands to: 6 + 8/8 + 2/8 Remember, 8/8 = 1 so 6 + 1 + 2/8 Becomes 7 + 2/8 Which reduces to 7 + 1/4

OpenStudy (anonymous):

I'm aware there are many ways to do this, but I chose the method he was probably taught to do.

hero (hero):

If you want to put that in improper fraction form it's 29/4

hero (hero):

Notice that 58/8 reduces to 29/4

hero (hero):

It's always good to one's benefit to know more than one way.

OpenStudy (anonymous):

I know, I was simply answering your question. @DannyFlash 1. A requires more as 3 3/8 > 3 1/4 2. Size 8 = 3 3/8, Size 14 = 3 7/8. 3 3/8 + 3 7/8 is what we just answered. 3. Similar question to (2), I'm sure you can work it out now. 4. Style A in size 10 requires 3 1/2 yards of fabric. What's half of this? 5. Style B in size 12 requires 3 7/8. How many times does this go into 12 1/8? The remainder is your answer.